OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. –2x + 3y = -6

OpenStudy (ivettef365):

first write the equation in form of y = mx + b m = slope b = y-int

OpenStudy (anonymous):

would you start at -6

OpenStudy (ivettef365):

no, first isolate the y on equation: -2x +3y = -6

OpenStudy (ivettef365):

do you know how to ?

OpenStudy (anonymous):

just put it by itself?

OpenStudy (ivettef365):

yes

OpenStudy (ivettef365):

do you know how to

OpenStudy (ivettef365):

after you do the y = mx + b you need to explain how to graph